### v2

# Focal Surface Holographic Light Transport using Learned Spatially Adaptive Convolutions

Abstract.
Computer-Generated Holography (CGH) is a set of algorithmic methods for identifying holograms that reconstruct Three-Dimensi-onal (3D) scenes in holographic displays. CGH algorithms decompose 3D scenes into multiplanes at different depth levels and rely on simulations of light that propagated from a source plane to a targeted plane. Thus, for $n$ planes, CGH typically optimizes holograms using $n$ plane-to-plane light transport simulations, leading to major time and computational demands. Our work replaces multiple planes with a focal surface and introduces a learned light transport model that could propagate a light field from a source plane to the focal surface in a single inference. Our learned light transport model leverages spatially adaptive convolution to achieve depth-varying propagation demanded by targeted focal surfaces. The proposed model reduces the hologram optimization process up to 1.5x, which contributes to hologram dataset generation and the training of future learned CGH models.
Computer-Generated Holography, Light Transport, Optimization, Spatially Adaptive Convolutions, Convolutional Neural Networks
copyright: ccccs: Hardware Emerging optical and photonic technologiesccs: Human-centered computing Displays and imagersccs: Computing methodologies Computer graphics
1. introduction
is a family of algorithmic methods used to generate holographic interference patterns. Identifying these interference patterns using learned (Shi et al., 2022) and optimization (Kavaklı et al., 2023a)  methods require conventional simulations of light propagation from plane-to-plane (Shen and Wang, 2006; Matsushima and Shimobaba, 2009). Recently, learned proxy methods (Choi et al., 2021; Kavaklı et al., 2022) have been proposed to replace conventional light propagation methods (Shen and Wang, 2006; Matsushima and Shimobaba, 2009). As these learned proxy methods for light propagation are trained using camera-in-the-loop strategies, they are able to capture imperfections of optical hardware, closing the gap between theoretical
Figure 1. Conventional Light Transport VS. Proposed Focal Surface Light Transport.(Source image: Tobi 87, Link: Wikimedia Commons)
simulations and actual hardware. Either learned or conventional, simulating light propagation among multiple planes in a 3D volume is computationally demanding, as a 3D volume is represented with multiple planes and each plane requires a separate calculation of light propagation to reconstruct the target image.
Our work introduces a learned focal surface light propagation model that could help free light simulations from plane dependence. Specifically, our model can propagate a phase-only hologram represented with a plane to a targeted focal surface, see Fig. 1. In our model, we extract  depth features of a focal surface by learning a set of  kernels. In addition, our model combines these  learned kernels with  kernels using a . Thus, effectively capturing  and  features of light propagation over a focal surface. Our work makes the following contributions:
  • Learned focal surface light transport model. By uniquely leveraging for , we introduce a new learned light transport model. Our model identifies a mapping from a phase-only hologram represented over a plane to a targeted focal surface.
  • Focal surface-based hologram optimization. To evaluate its practicality, we utilize our model for a 3D phase-only hologram optimization application. Compared with conventional light propagation based hologram optimization methods (Kavaklı et al., 2023b, a), our approach accelerates the optimization process up to 1.5x, leading to speed up benefits in hologram dataset generation and training future learned  models.
  • Experimental Validation. We evaluate our method in simulation for various propagation distances and validate the result using a bench-top on-axis holographic display prototype.
Figure 2. Our proposed learned focal surface light transport model. The process starts with an input hologram $\mathbf{H}$ and a focal surface $\mathbf{D}$ to generate spatially varying kernels $[\mathbf{V_{i}}]$, where $i=0,1,2,3$ indicates the index of scales. Those kernels are utilized in the Spatially Adaptive Module (SAM) to achieve focal surface light transport. In the SAM, $\mathbf{V}_{3}^{0},\mathbf{V}_{3}^{j},\mathbf{V}_{3}^{l},\mathbf{V}_{3}^{z}$ represent kernels used at different spatial locations, where 00, $j$, $l$, and $z$ indicate specific positions. (Source image: Tobi 87, Link: Wikimedia Commons)
2. Focal Surface Light Transport
We introduce the , a modified convolution structure for encoding  features. Leveraging the , our work enables the learned focal surface light transport network.
2.1. Spatially Adaptive Convolution
Standard Convolution
Given an input feature $\mathbf{\tilde{I}}\in\mathbb{R}^{\tilde{c}\times\tilde{h}\times\tilde{w}}$ in a , where $\tilde{c}$, $\tilde{h}$, and $\tilde{w}$ represent the number of channels, height, and width of the input $\mathbf{\tilde{I}}$ (in our case, $\tilde{c}=3$, $\tilde{w}=1080$, $\tilde{h}=1920$), the discrete convolution based on a  kernel $\mathbf{W}\in\mathbb{R}^{\hat{c}\times\tilde{c}\times k\times k}$ is defined as:
(1) $\underbrace{\mathbf{I}[c,x,y]}_{{\text{output}}}=\sum_{c^{\prime},x^{\prime},y^{\prime}}\underbrace{\mathbf{W}[c,c^{\prime},x^{\prime},y^{\prime}]}_{{\text{Kernel}}}\underbrace{\mathbf{\tilde{I}}[c^{\prime},x+x^{\prime},y+y^{\prime}]}_{{\text{input}}},$
where $\tilde{c}$ and $\hat{c}$ indicate the number of input and output channels. The indices satisfy $1\leq c^{\prime}\leq\tilde{c}$ and $1\leq c\leq\hat{c}$. The pair $(x^{\prime},y^{\prime})$ belongs to the set $\Omega(k)$, which specifies a $k\times k$ convolutional window. The summation operation acts on all input channels, which implies that each input channel contributes to every output channel. According to Eq. (1), this operation is characterized by a kernel that is spatially shared and content-independent. Learning-based light transport models could use Eq. (1) as a basic operation. However, it is challenging for this method to project a hologram onto a focal surface. As each pixel on the hologram plane may correspond to a different depth on the focal surface, which makes the  kernel a sub-optimal choice to capture  features (Zheng et al., 2021; Xu et al., 2020), including focusing or out-of-focus effects due to depth variance. A typical solution is to employ a large number of parameters for feature encoding, resulting in an increased memory footprint. Alternatively, we could consider using  convolution (Zheng et al., 2021; Xu et al., 2020). The  kernel $\mathbf{V}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ incorporates two new dimensions $h,w$ into  kernel, where $h$ and $w$ indicate height, and width of the output feature. However, relying solely on  kernels may increase model capacity due to extra parameters, particularly when $h$ and $w$ are large. These alternative designs all demand extra network capacity.
Spatially Adaptive Convolution Operation
To address these problems, we utilize the  based on (Xu et al., 2020). Our method reduces the network parameters by multiplying the  kernel with the standard  kernel. Initially, the  kernel $\mathbf{V}\in\mathbb{R}^{1\times h\times w\times\tilde{c}\times k\times k}$ is introduced, the output channel is set to $1$ to reduce the number of parameters. The  kernel $\mathbf{A}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ is computed by multiplying the $\mathbf{W}$ and $\mathbf{V}$, which defined as:
(2) $\mathbf{A}[c,x,y,c^{\prime},x^{\prime},y^{\prime}]=\mathbf{V}[1,x,y,c^{\prime},x^{\prime},y^{\prime}]*\mathbf{W}[c,c^{\prime},x^{\prime},y^{\prime}],$
where $1\leq c\leq\hat{c}$, $1\leq c^{\prime}\leq\tilde{c}$ ,$1\leq x\leq h$ and $1\leq y\leq w$. Eq. 2 enhances the output channel capacity in $\mathbf{V}$ while maintaining spatially variant. Both $\mathbf{V}$ and $\mathbf{W}$ can be either pre-defined or learned, making the network content-adaptive. By using $\mathbf{A}$, the  is defined as:
(3) $\mathbf{I}[c,x,y]=\sum_{c^{\prime},x^{\prime},y^{\prime}}\underbrace{\mathbf{A}[c,x,y,c^{\prime},x^{\prime},y^{\prime}]}_{{\text{SA Kernel}}}\mathbf{\tilde{I}}[c^{\prime},x+x^{\prime},y+y^{\prime}].$
retains both the dimensional coherence of the  kernel in  and is spatially variant at the same time. Note that when $\mathbf{W}$ becomes an all-one tensor, Eq. 3 is equivalent to the  convolution in .
[…]
Ground Truth
[…]
Ours
[…]
U-Net
[…]
In focus
             Out of focus
                         In focus
[…]
In focus
             Out of focus
                         In focus
Figure 3. Visual comparison of simulating light transported onto a focal surface (specified in the first row of each case) at 0 mm and 10 mm propagation distances. The ground truth is obtained via ASM (Matsushima and Shimobaba, 2009). Both focused and defocused regions indicate poor performance of the U-Net model. (Source image: Matt H. Wade, Link: Wikimedia Commons)
2.2. Learned Focal Surface Light Transport
We first generate  kernels to encode depth-varying features of the focal surface, which are later used in  for focal surface light transport. For the schematic figure of our system, please see Fig. 2.
Spatially Varying Kernel Generation
As shown in Fig. 2, the  kernel generation module takes the hologram $\mathbf{H}\in\mathbb{R}^{1\times 3\times h\times w}$ and focal surface $\mathbf{D}\in\mathbb{R}^{1\times 1\times h\times w}$ as inputs. We adopted the architecture in RSGUNet (Huang et al., 2018) for  kernel generation module. The output of each decoder layer is integrated with feature maps from different layers in the encoders. Then combined features will be fed into  module to learn a set of  kernels $[\mathbf{V}_{i}]$, where $\mathbf{V}_{i}\in\mathbb{R}^{n\times\tilde{c}_{i}\times k\times k}$, $i=0,1,2,3$ refers to different scale levels, $\tilde{c}_{i}$ denotes the input channel, $k$ is the kernel size, and $n=\frac{h}{2^{i}}\times\frac{w}{2^{i}}$ is the number of kernels. The  module contains convolution layers and average pooling layers. To mitigate artifacts, we modify the global feature module in (Huang et al., 2018) to an attention block and apply it at the bottleneck of the U-Net.
Focal Surface Light Transport
We leverage the generated kernels to build our light transport module based on RSGUNet (Huang et al., 2018). The module takes the hologram $\mathbf{H}$ as input without requiring depth, as the depth feature of the focal surface is inherently encoded within the learned  kernels. To integrate the  features into the encoder, we propose a  based on . As shown in Fig 2, we first replace the  kernel $\mathbf{W}$ to an all-ones tensor in Eq. (2), which ignores the   kernels and only considers the  kernels to capture the original  information. In parallel, we introduce the  kernels back to Eq. (2) as a learning parameter and multiply with the  kernels for better diverse feature extraction. These features from the two operations will be concatenated to form the output of . Finally, the global feature module and the decoder will process the output to generate the reconstruction at the given focal surface denoted as $\mathbf{R}$.
Loss function
We employ the $L_{2}$ norm to quantify the discrepancy between the reconstruction $\mathbf{R}$ and the target image $\mathbf{R}^{\prime}$. Both $\mathbf{R}$ and $\mathbf{R}^{\prime}$ are focal surface depended image reconstructions. Since $\mathbf{R}$ contains both focus and defocus regions (Kavaklı et al., 2023a), we utilize a binary mask $\mathbf{M}$ that highlights only the focus parts of the image. The loss function for the reconstruction on a single focal surface $\mathcal{L}_{D}$ is defined as:
(4) $\mathcal{L}_{D}=\alpha_{0}\mathbf{M}\|\mathbf{R}-\mathbf{R}^{\prime}\|_{2}^{2}+\alpha_{1}(\mathbf{1}-\mathbf{M})\|\mathbf{R}-\mathbf{R}^{\prime}\|_{2}^{2},$
where $\alpha_{0}$ and $\alpha_{1}$ represent weights ($\alpha_{0}=1$ and $\alpha_{1}=0.5$).
2.3. Optimizing Holograms with Focal Surfaces
Recently, learning-based methods have been proposed to solve 3D hologram generation tasks (Choi et al., 2021; Shi et al., 2022). However, the ideal 3D hologram for the holographic display has not yet been precisely defined (Kim et al., 2024). Optimization-based hologram generation methods (Kavaklı et al., 2023b, a) could potentially help identify the ideal 3D hologram and generate hologram datasets for learning-based approaches. Typically, optimization methods are based on the multiplane representation, where a full-color hologram is synthesized by making use of the phase patterns of the three color primaries. Following previous work (Kavaklı et al., 2023b), each single-color phase pattern is obtained by:
(1) $\mathbf{\hat{H}}_{p}\leftarrow\arg\min_{\mathbf{H_{p}}}\sum_{p=1}^{3}\mathcal{L}\left(\left|e^{i\mathbf{H}_{p}}\otimes\mathbf{K}_{p}\right|^{2},s\mathbf{R}_{p}\right),$
where $p$ denotes the index of a color primary, $\mathbf{H}_{p}$ is the SLM phase, $\mathbf{\hat{H}}_{p}$ is the optimized SLM phase, $\mathbf{K}_{p}$ is the wavelength-dependent light transport kernel (Matsushima and Shimobaba, 2009), $\mathbf{R}_{p}$ is the target image intensity, $s$ is an intensity scaling factor ( $s=1$ by default), $\otimes$ denotes convolution. We substitute the conventional light transport model with our focal-surface-based model:
(5) $\mathbf{\hat{H}}\leftarrow\arg\min_{\mathbf{H}}\mathcal{L}\left(F(\mathbf{\mathbf{H},\mathbf{D}}),s\mathbf{R}\right).$
In this case, the hologram optimization problem is simplified. Our approach simultaneously optimizes hologram in three color primaries and maintains phase-only at the same time.
3. Evaluation and Discussion
[…]
Rear focus
 Front focus
  Rear focus
  Front focus
Out of focus
Out of focus
In focus
In focus
In focus
Out of focus
Out of focus
Out of focus
In focus
In focus
In focus
Out of focus
Figure 4. Visual comparison on simulated holograms optimized using ASM 6 and Ours 6 under 0 mm propagation distance. All holograms are reconstructed using ASM for evaluation. (Source image : Jaimie Phillips, Link: Wikimedia Commons)
We generate the focal surface light transport dataset based on previous work (Kavaklı et al., 2023b, a) at the resolution $1920\times 1080$. See Section 1 of the supplementary material for more details. We use Adam optimizer $(\beta_{1}=0.9,\beta_{2}=0.999,\alpha_{decay}=0.5~{}after~{}50~{}epochs)$. The model is trained for $500$ epochs, with an initial  of $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA V100 16G GPU.
Evaluation.
To assess the image quality, we utilize metrics including Peak Signal-to-noise Ratio (PSNR), Structural Similarity (SSIM), and Perceptual Similarity Metric (LPIPS) (Zhang et al., 2018). First, we assess the quality of light simulation on a focal surface. As shown in Tbl. 1, our model outperforms U-Net (Ronneberger et al., 2015) across all metrics. Fig. 3 shows that our model preserves more high-frequency content than U-Net, providing finer details and sharper edges, closer to the ground truth.
Table 1. Evaluation of various light transport models on our dataset. The speed is tested by simulating an all-in-focus, full-color 3D image with six depth planes. Note that higher PSNR/SSIM and lower Params/Speed indicate better performance, denoted by $\uparrow$ and $\downarrow$ in the tables.
Methods PSNR (dB) $\uparrow$ SSIM $\uparrow$ Stage Params $\downarrow$ Speed $\downarrow$
0 mm/10 mm 0 mm/10 mm (M) (s)
ASM (GT) (Matsushima and Shimobaba, 2009) - - Two - 0.4559
U-Net (Ronneberger et al., 2015) 29.662/30.112 0.8015/0.7760 Single 7.7760 0.0565
Ours $\mathbf{36.016}$/$\mathbf{34.279}$ $\mathbf{0.9128}$/$\mathbf{0.8470}$ Single $\mathbf{7.4446}$ $\mathbf{0.0471}$
Second, we utilize our model for a 3D phase-only hologram optimization application under $0mm$ propagation distance. Optimizing holograms with six target planes using  (Matsushima and Shimobaba, 2009) is denoted as ASM 6, while Ours 4 and Ours 6 represent optimizing holograms using our model with four and six focal surfaces, respectively.
Table 2. Comparison of image quality for the scene in Fig. 4 among ASM 6, Ours 6, and Ours 4 across different iterations at 0 mm propagation distance. Note that higher PSNR/SSIM $\uparrow$ and lower LPIPS/Speed $\downarrow$ indicate better performance.
ASM 6///Ours 6 /// Ours 4 Iteration
50 100 200
Speed (s) $\downarrow$ 42.580/30.182/$\mathbf{20.869}$ 84.626/61.460/$\mathbf{39.792}$ 171.49/119.02/$\mathbf{77.878}$
PSNR (dB) $\uparrow$ 27.377/$\mathbf{27.501}$/26.088 $\mathbf{27.795}$/27.598/26.905 $\mathbf{27.801}$/27.625/26.928
SSIM $\uparrow$ $\mathbf{0.7100}$/0.6868/0.6142 $\mathbf{0.7193}$/0.6933/0.6753 $\mathbf{0.7195}$/ 0.6890/0.6767
LPIPS $\downarrow$ $\mathbf{0.3971}$/0.4747/0.5431 $\mathbf{0.3894}$/0.4687/0.4707 $\mathbf{0.3889}$/0.4787/0.4689
All holograms are reconstructed using ASM for performance assessment. As shown in Fig. 4 and Tbl. 2,  Ours 6 achieves comparable results with about $70\%$ of the optimization time compared to ASM 6. Actual captures of Ours 6 and ASM 6 in Fig. 5 demonstrate the capability of our model for generating 3D holograms. For more details on the display prototype and comparisons, see Sections 2 and 3 in supplementary material.
  Ours 6
  ASM 6
Figure 5. Comparing experimental captures of  ASM 6 and Ours 6 under 0 mm propagation distances. (Source image : Jaimie Phillips, Link: Wikimedia Commons)
Computational Complexity Analysis.
First, we assess the computational complexity of simulating a full-color, all-in-focus 3D image across six depth planes. As shown in Tbl. 1, conventional ASM-based model (Matsushima and Shimobaba, 2009) requires eighteen forward passes to simulate a full-color, all-in-focus 3D image with six depth planes. In contrast, our model simulates the three color-primary images simultaneously onto a focal surface with a single forward pass, reducing simulation time by 10x and achieving better image quality with fewer parameters compared to U-Net (Ronneberger et al., 2015). Second, we evaluate hologram optimization. In Tbl. 2, using four focal surfaces (Ours 4) to approximate six planes for focus and defocus guidance, speeding up optimization by up to 2x. Increasing the number of focal surfaces to six (Ours 6) achieves comparable results with about a 1.5x speedup.
Limitations and Future Works
As shown in Fig. 3, the performance of our model degrades at a long propagation distance $(10\ mm)$ compared to zero distance $(0\ mm)$. See Section 3 in the supplementary material for more comparisons. Future improvements could include using a factorized larger kernel for long-distance propagation. In addition, our model focuses on depth-varying propagation within a 3D volume, more investigation is needed for depth-varying propagation of the entire volume using conditional networks.
Acknowledgements.
The authors thank reviewers for their valuable feedback; Louise L. Xie for her feedback on the manuscript; Ziyang Chen and Doğa Yılmaz for their discussions and assistance. We also thank Northeastern University for computing resources for early experiments.